# In Context Learning and Reasoning for Symbolic Regression with Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) are transformer-based machine learning models that have shown remarkable performance in tasks for which they were not explicitly trained. Here, we explore the potential of LLMs to perform symbolic regression — a machine-learning method for finding simple and accurate equations from datasets. We prompt GPT-4 and GPT-4o models to suggest expressions from data, which are then optimized and evaluated using external Python tools. These results are fed back to the models, which proposes improved expressions while optimizing for complexity and loss. Using chain-of-thought prompting, we instruct the models to analyze data, prior expressions, and the scientific context (expressed in natural language) for each problem before generating new expressions. We evaluated the workflow in rediscovery of five well-known scientific equations from experimental data, and on an additional dataset without a known equation. GPT-4 and GPT-4o successfully rediscovered all five equations and generally performed better when prompted to use a scratchpad and consider scientific context. GPT-4o model demonstrated improved reasoning with data patterns, particularly evident in the dual-site Langmuir and Nikuradse datasets. We demonstrate how strategic prompting improves the model’s performance and how the natural language interface simplifies integrating theory with data. We also observe how theory can sometimes offset noisy data and, in other cases, data can make up for poor context. Although this approach does not outperform established SR programs where target equations are more complex, LLMs can nonetheless iterate toward improved solutions while following instructions and incorporating scientific context in natural language.
@@ 1 Introduction
Data analysis is ubiquitous in all disciplines, where identifying correlations between variables is key to finding insights, informing conclusions, supporting hypotheses, or developing a new theory. For scientific data, we often aim to find expressions with few adjustable parameters explaining the data while ensuring that they align with theory. Symbolic regression is a machine learning technique that approaches equation-based scientific discovery – given a dataset, it searches through some “space of possible equations” and identifies those that balance accuracy and simplicity. It is different from conventional regression methods, as symbolic regression infers the model structure from data rather than having a predetermined model structure.
@@
Mathematically, symbolic regression is formulated as some form of optimization, not just of the constants in an equation, but as a search through “equation space" for optimal expressions. In this way, symbolic regression is a form of machine learning – as data is received, an internal model is updated to match the data; when the model fits the data well and can make predictions about unseen data, the algorithm is said to have “learned" the underlying patterns in the data. In contrast to popular machine learning algorithms like neural networks, symbolic regression does not just only fit the constants in an equation but also finds functional forms that match the data.
@@
Symbolic regression methods mainly use genetic algorithms [1, 2, 3, 4] that generate random expressions from data, optimize their parameters, and evaluate their fitness with respect to the data through an iterative process until a fitness level or a specific number of iterations is reached. Other approaches include using Markov chain Monte Carlo (MCMC) sampling [5, 6], mixed integer nonlinear programming [7, 8, 9], greedy tree searches [10], pre-trained transformer-based models [11, 12], and sparse matrix algorithms [13, 14, 15]. These techniques are broadly geared towards accelerating equation search or efficient multi-objective optimization, but they do not integrate reasoning. Researchers have long explored ways to make these algorithms more informed by guiding the search space based on the context of the data [16, 17, 18, 19, 6, 20, 21, 22, 23, 24, 25, 26, 27, 9, 28], including using large language models integrated with genetic algorithms [29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. As scientific data is strongly tied to theory, encoding them in the program narrows the vast search space and can make the programs more effective.
@@
This work explores an approach to symbolic regression using large language models (LLMs) for equation discovery. LLMs are machine learning models adept at understanding and generating natural language. At its core, an LLM uses the transformer architecture—a neural network developed by Google that scales very effectively and allows the training of models on massive datasets [39]. The term “large” in the language model refers to the size and intricacy of the network, along with the dataset on which it was trained. Prior to GPT-3, natural language processing (NLP) tasks were solved by pretraining language models on vast text datasets and fine-tuning them for specific tasks. However, GPT-3 demonstrated that language models can excel at tasks using in-context learning without necessitating fine-tuning [40, 41]. LLMs are now commonly used for tasks like chat, code generation, summarization, translation, etc. -- and quite remarkably, these tasks can be effectively accomplished by using English language as model input without the need for machine learning expertise. We have firsthand experience with these models’ capabilities by interacting with AI chatbots on platforms like ChatGPT, Claude, or Gemini. LLMs have a wide understanding of the world from their training data and can even solve simple math problems expressed in natural language [42, 43, 44]. They are contributing significantly in education and research [45, 46, 47, 48] medicine [49, 50], physical and social sciences [51, 52, 53], as well as in legal [54, 55], business [56, 57], and entertainment [58, 59, 60, 61] sectors.
@@
Existing transformer-based symbolic regression programs [62, 11, 12] use models pretrained on large databases of synthetically-generated dataset/expression pairs, designed specifically for symbolic regression tasks. Thus, these approaches learn to pattern-match between datasets and math expressions, but they don’t employ iterations like those in genetic programming that optimize expressions for complexity and loss. In that respect, LLMs have been used to imitate evolutionary algorithms (EA) [29, 30, 31, 32, 33, 34, 35], and have specifically been applied to solve symbolic regression (SR) problems [31, 34]. Meyerson and coworkers [31] developed a workflow that performs genetic programming (mutation, crossover, etc.) through prompts in LLMs and tested symbolic regression in two ways: first, by using a language model in all evolutionary operators except the fitness measure, and second, by only using a language model in the initialization, crossover, and mutation operators. The second approach more closely resembles our work, but what we do is even simpler: we task LLMs to generate and/or transform expressions freely. Furthermore, while these approaches leverage “in-context” learning and don’t require pretraining [31, 34], the context employed is limited to a list of previously-obtained expressions. We propose to expand the context to include data, as well natural language descriptions of the scientific context of the problem. To effectively use this context, we anticipate that the LLM will perform better if given time for analysis [63, 64, 65]. Therefore, we also incorporate zero-shot chain-of-thought prompting with a scratchpad [66] to frame equation generation for symbolic regression as a reasoning problem [67] in the context of the data and free-form scientific information.
@@
Symbolic regression requires equation generation and precise fitting of numerical constants. Yang and coworkers [68] show that LLMs can perform linear regression, optimizing constants in math expressions via feedback loops, without showing the LLM the analytical form. They find that “LLM can often calculate the solution directly from the analytic form," however we find LLMs to be unreliable and inefficient for such tasks. Consequently, we interleave LLM-based optimization with gradient-based optimization, following a similar approach as [33], iteratively refining prompts for more accurate output. In our work, an LLM guides optimization of the symbolic structure of the math expressions, while SciPy performs numerical optimization of the constants. This approach aligns closely with two recent works [36, 37] in the literature where LLMs were used to generate expressions and optimizations were performed via iterations. Although there are some variations in the workflow and significant differences in the explored datasets, the unique contribution of this work lies in using both data and context in natural language as model input, and implementing a scratchpad to record raw model output as it solves the problem. This promotes informal reasoning about the data and context, and helps reveal “test set leakage.”
@@ 2 Methods
In our initial tests, we asked LLMs to generate mathematical expressions from scientific data. We used a simple prompt to assess its capability and tested GPT-3.5-turbo and GPT-4 at varying temperatures (Figures 1). In all cases, the models produced expressions while “hallucinating” arbitrary coefficients. We revised the prompts to ask the LLMs to “show all steps" to gain insight into how the model selects these values [65]. In response, the output either provided Python code for optimization or, at times, mathematical steps for optimization.
@@
Figure 1: Illustration of GPT-4 attempting symbolic regression. GPT-4 predicts expressions with optimized coefficients when passed a dataset for nitrogen adsorption on mica [69]. The Python code snippet from Prompt 2 output has been truncated to keep the figure concise. Note that the actual parameter values produced by running the code differ from what GPT-4 generates. Figure 16 in SI shows the output from GPT-3.5 turbo which produces similar results.
@@
Although the generated code accurately performed curve-fitting, the LLM hallucinated incorrect coefficients. Nonetheless, the generated remarks about the data patterns hinted that they may have the potential to generate accurate functional forms that can be optimized outside the LLM.
System design
@@
Therefore, we designed a workflow (Figure 2) where we task GPT-4 and GPT-4o with suggesting expressions without fitting constants, and subsequently, we optimize the coefficients of the expressions using SciPy outside the LLM. A Python class takes in expressions, optimizes them, then calculates their complexity and mean squared error (MSE). These results are stored as a dictionary, the text of which is passed back to LLMs in a subsequent prompt asking to suggest better expressions. We initially evaluated GPT-3.5-turbo, but found it less reliable in following instructions than GPT-4 and GPT-4o, more frequently generating expressions that didn’t parse.
@@
Figure 2: Workflow for using LLMs for SR. First, the dataset is sent to LLMs (with or without context) which is instructed to suggest expressions without optimizing parameters. The generated expressions are then parsed by Python and optimized using SciPy (Nelder-Mead [70] method with basin-hopping as the numerical optimizer [71]). Results for each expression are stored in a Python dictionary, and added to a list of results from prior expressions. The top performing expressions are sent as feedback to LLMs, which is asked to suggest better expressions optimizing for both complexity and loss. The feedback loop is run for a set number of iterations.
@@
We use two prompts: 1) Initial Prompt - where we input data and ask GPT-4 and GPT-4o to suggest expressions, and 2) Iteration Prompt - where, along with data, we provide feedback in JSON format. We sort the expressions in descending order of mean squared error, then share this list as feedback, asking the LLMs to suggest new equations optimizing for complexity and loss. In addition to this, we also include a system message to guide the behavior of the language model, setting the tone of the conversation. We do not use chat history or any advanced forms of memory [72, 73, 74, 75] in this workflow. The history of generated expressions is maintained externally and provided as feedback in the iteration prompt. A Python function maintains this feedback loop by sorting and filtering a list of dictionaries based on MSE and complexity. At the start of the search, up to 6 expressions are always returned; later, the least accurate expressions that are not on the Pareto front are pruned in order to manage the length of the context window.
By making each call to LLMs independent, we aim to minimize hallucinations that have been observed when the chat history becomes too large in conversational models [76, 77], as well as manage cost by pruning the large quantity of generated scratchpad text.
@@
Removing bias: We aimed to make the workflow run smoothly without any human intervention and therefore, it was important to obtain machine-readable and precise output from LLMs to ensure the SciPy function runs without any errors. A simple way to illustrate the expected outcome was by providing examples in a few-shot context [40, 82]. While this led to expressions matching the required syntax, we noticed the generated expressions resembled the examples we provided (Figure 3). While this taught the LLM correct syntax, it introduced bias that severely compromised the search.
Figure 3: Bias from examples GPT-4 generated exact expressions from the prompt examples that were given to illustrate the output syntax. The revised prompts (Prompt 2 and 3) eliminate this bias from the equation prediction process.
This motivated our two-prompt setup, with an initial prompt tasking LLMs to generate unbiased expressions in LaTeX and an iteration prompt receiving previously-generated examples as Python script (now without bias). To bridge these, we converted the LaTeX text into SciPy-compatible text using a Python function for string formatting, which we developed after identifying the types of expressions GPT-4 was prone to generating. Although this approach did not completely resolve syntax errors, it effectively managed them for the GPT-4 model. While running the GPT-4o model, we realized that most syntax errors stemmed from incorrect LaTeX formatting. To address this, we made a minor adjustment by removing the term "LaTeX" and allowing the model to generate expressions as Python strings instead. This small change was was sufficient to remove bias and was also less prone to syntax errors. We nervetheless processed the generated Python expressions through our formatting function, and it worked seamlessly.
@@
(a) Effect of scratchpad on Langmuir Dataset
(b) Effect of scratchpad on Kepler Law Dataset
Figure 4: Illustration of scratchpad approach. We observe substantial, qualitative improvements in the predicted expressions after implementing the scratchpad technique (Prompt 2). The suggested expressions for (a) Langmuir’s and (b) Kepler’s Law dataset include operators (/// and sqrt, respectively) present in the target models ($y=\frac{c_{1}*x}{c_{2}+x}$ and $y^{2}=c_{1}x^{\frac{3}{2}}$, respectively)
@@
Consider scientific context: Our primary motivation for building this system was to test whether providing scientific context could shape the expressions generated by the LLM. SR programs often successfully generate expressions that fit the data well and are simple, yet they may not adhere to scientific principles or be otherwise “meaningful.” Yet scientists often have valuable insights into their domain that extend beyond these constraints, and they may not always know which specific expressions will best capture the nuances of their dataset or if entirely new expressions might be more effective. By incorporating scientific context, we aim to align the equation search to be consistent with scientific theories.
Classical SR approaches always incorporate some amount of guidance from the practitioners (e.g. by limiting available math operators and variables and incorporating some bias toward parsimonious expressions); more recent work enables SR programs to explicitly account for limiting behavior (shape-constrained SR) and dimensional constraints [84, 85, 86, 24, 18, 25, 87]. Incorporating these into SR algorithms typically requires bespoke modification of research software; our strategy is to use LLM prompting in natural language to instill this context, which might additionally include more fluid constraints like “generate diverse equations" and “consider scientific context." For instance, the context we provide for Kepler’s Law is a single line text stating, “The data is about planetary motion in astrophysics where the independent variable (x1) is semi-major axis, and the dependent variable (y) is period in days." Figure 17 in SI lists the context provided for each of the datasets.

### v3

@@ Abstract
Large Language Models (LLMs) are transformer-based machine learning models that have shown remarkable performance in tasks for which they were not explicitly trained. Here, we explore the potential of LLMs to perform symbolic regression – a machine-learning method for finding simple and accurate equations from datasets. We prompt GPT-4 and GPT-4o models to suggest expressions from data, which are then optimized and evaluated using external Python tools. These results are fed back to the LLMs, which propose improved expressions while optimizing for complexity and loss. Using chain-of-thought prompting, we instruct the models to analyze data, prior expressions, and the scientific context (expressed in natural language) for each problem before generating new expressions. We evaluated the workflow in rediscovery of Langmuir and dual-site Langmuir’s model for adsorption, along with Nikuradse’s dataset on flow in rough pipes, which does not have a known target model equation. Both the GPT-4 and GPT-4o models successfully rediscovered equations, with better performance when using a scratchpad and considering scientific context. GPT-4o model demonstrated improved reasoning with data patterns, particularly evident in the dual-site Langmuir and Nikuradse dataset. We demonstrate how strategic prompting improves the model’s performance and how the natural language interface simplifies integrating theory with data. We also applied symbolic mathematical constraints based on the background knowledge of data via prompts and found that LLMs generate meaningful equations more frequently. Although this approach does not outperform established SR programs where target equations are more complex, LLMs can nonetheless iterate toward improved solutions while following instructions and incorporating scientific context in natural language.
@@ 1 Introduction
Data analysis is ubiquitous in all disciplines, where identifying correlations between variables is key to finding insights, informing conclusions, supporting hypotheses, or developing a new theory. For scientific data, we often aim to find expressions with few adjustable parameters explaining the data while ensuring that they align with theory. Symbolic regression is a machine learning technique that approaches equation-based scientific discovery. Given a dataset, it searches through some “space of possible equations" and identifies those that balance accuracy and simplicity. It is different from conventional regression methods, as symbolic regression infers the model structure from data rather than having a predetermined model structure.
@@
Mathematically, symbolic regression is formulated as some form of optimization, not just of the constants in an equation, but as a search through “equation space” for optimal expressions. In this way, symbolic regression is a form of machine learning – as data is received, an internal model is updated to match the data. When the model fits the data well and can make predictions about unseen data, the algorithm is said to have “learned” the underlying patterns in the data. In contrast to popular machine learning algorithms like neural networks, symbolic regression not only fits the constants in an equation but also finds functional forms that match the data.
@@
Symbolic regression methods mainly use genetic algorithms [59, 54, 86, 66] that generate random expressions from data, optimize their parameters, and evaluate their fitness with respect to the data through an iterative process until a fitness level or a specific number of iterations is reached. Other approaches include using Markov chain Monte Carlo (MCMC) sampling [49, 40], mixed integer nonlinear programming [5, 24, 23], greedy tree searches [27], pre-trained transformer-based models [50, 26], and sparse matrix algorithms [15, 71, 80]. These techniques are broadly geared towards accelerating equation search or efficient multi-objective optimization, but they do not integrate reasoning. Researchers have long explored ways to make these algorithms more informed by guiding the search space based on the context of the data [36, 70, 68, 93, 40, 16, 56, 30, 55, 42, 92, 53, 72, 23, 33], including using large language models integrated with genetic algorithms [67, 61, 75, 11, 41, 43, 65, 73, 89, 38]. As scientific data is strongly tied to theory, encoding it in the program narrows the vast search space and can make the programs more effective.
@@
This work explores an approach to symbolic regression using large language models (LLMs) for equation discovery. LLMs are machine learning models adept at understanding and generating natural language. At its core, an LLM uses the transformer architecture—a neural network developed by Google that scales very effectively and allows the training of models on massive datasets [96]. The term “large” in the language model refers to the size and intricacy of the network, along with the dataset on which it was trained. Prior to GPT-3, natural language processing (NLP) tasks were solved by pretraining language models on vast text datasets and fine-tuning them for specific tasks. However, GPT-3 demonstrated that language models can excel at tasks using in-context learning without necessitating fine-tuning [14, 9]. LLMs are now commonly used for tasks like chat, code generation, summarization, translation, etc. - and quite remarkably, these tasks can be effectively accomplished by using English language as model input without the need for machine learning expertise. We have firsthand experience with these models’ capabilities by interacting with AI chatbots on platforms like ChatGPT, Claude, or Gemini. LLMs have a wide understanding of the world from their training data and can even solve simple math problems expressed in natural language [94, 44, 84]. They are contributing significantly in education and research [51, 57, 35, 32] medicine [22, 69], physical and social sciences [31, 87, 82], as well as in legal [17, 39], business [8, 37], and entertainment [46, 20, 90, 45] sectors.
@@
Existing transformer-based symbolic regression programs [95, 50, 26] use models pretrained on large databases of synthetically-generated dataset/expression pairs, designed specifically for symbolic regression tasks. Thus, these approaches learn to pattern-match between datasets and math expressions, but they don’t employ iterations like those in genetic programming that optimize expressions for complexity and loss. In that respect, LLMs have been used to imitate evolutionary algorithms (EA) [67, 61, 75, 11, 41, 43, 65], and have specifically been applied to solve symbolic regression (SR) problems [75, 43]. Meyerson and coworkers [75] developed a workflow that performs genetic programming (mutation, crossover, etc.) through prompts in LLMs and tested symbolic regression in two ways: first, by using a language model in all evolutionary operators except the fitness measure, and second, by only using a language model in the initialization, crossover, and mutation operators. The second approach more closely resembles our work, but what we do is even simpler: we task LLMs to generate and/or transform expressions freely. Furthermore, while these approaches leverage “in-context” learning and don’t require pretraining [75, 43], the context employed is limited to a list of previously-obtained expressions. We propose to expand the context to include data, as well as natural language descriptions of the scientific context of the problem. To effectively use this context, we anticipate that the LLM will perform better if given time for analysis [6, 18, 98]. Therefore, we also incorporate zero-shot chain-of-thought prompting with a scratchpad [79] to frame equation generation for symbolic regression as a reasoning problem [19] in the context of the data and free-form scientific information.
@@
Symbolic regression requires equation generation and precise fitting of numerical constants. Yang and coworkers [101] show that LLMs can perform linear regression, optimizing constants in math expressions via feedback loops, without revealing the LLM the analytical form. They find that ‘LLM can often calculate the solution directly from the analytic form,’ however, we find LLMs to be unreliable and inefficient for such tasks [34]. Consequently, we interleave LLM-based optimization with gradient-based optimization, following a similar approach as [41], iteratively refining prompts for more accurate output. In our work, an LLM guides optimization of the symbolic structure of the math expressions, while SciPy performs numerical optimization of the constants. This approach aligns closely with two recent works [73, 89] in the literature where LLMs were used to generate expressions and optimizations were performed via iterations. Although there are some variations in the workflow and significant differences in the explored datasets, the unique contribution of this work lies in using data, context, and theoretical constraints in natural language as model input and implementing a scratchpad to record raw model output, as it solves the problem. This promotes informal reasoning about the data and context, and helps reveal “test set leakage.”
@@ 2.1 Preliminary tests
In our initial tests, we asked LLMs to generate mathematical expressions from scientific data. We used a simple prompt to assess its capability and tested GPT-3.5-turbo and GPT-4 at varying temperatures. In all cases, the models produced expressions while “hallucinating” arbitrary coefficients. We revised the prompts to ask the LLMs to “show all steps” to gain insight into how the model selects these values [98]. In response, the output either provided Python code for optimization or at times, mathematical steps for optimization, but with the wrong output.
@@
Figure 1: GPT-4’s hallucinated curve fitting on adsorption data. The plot shows how GPT-4 predicts expressions with coefficients when passed a dataset for nitrogen adsorption on mica [60]. The Python code that it generates is correct; however, the actual optimized parameter values produced by running the code differ from what GPT-4 outputs. Figure 19 and 20 in SI show the raw output from GPT-3.5 turbo and GPT-4.
@@
As illustrated in Figure 1, although the LLM-generated code runs and accurately performs curve-fitting, the model still hallucinates incorrect coefficients as its final output. Nonetheless, the generated remarks about the data patterns suggested that they may have the potential to generate accurate functional forms that can be optimized outside the LLMs.
2.2 System design
@@
Therefore, we designed a workflow (Figure 2) where we task models GPT-4 and GPT-4o with suggesting expressions without fitting constants, and subsequently, we optimize the coefficients of the expressions using SciPy outside the LLM. A Python class takes in expressions, optimizes them, and then calculates their complexity and mean squared error (MSE). The results are then stored as a dictionary, the text of which is passed back to LLMs in a subsequent prompt asking to suggest better expressions. We initially evaluated GPT-3.5-turbo, but found it less reliable in following instructions than GPT-4 and GPT-4o, more frequently generating expressions that didn’t parse. Although LLMs could have been used to generate the curve-fitting code, we ensure reliability and prevent reward hacking by writing and validating the code ourselves.
We define complexity as the number of operators and variables/constants in a given expression, or in other words, the total nodes in an expression tree (see Figure 3). A mathematical expression can be represented as a rooted tree structure where internal nodes correspond to operators, which can be binary (e.g., ++, *, //, or ˆ) or unary (like sin or log). The leaves, or terminal nodes, represent either variables (e.g., $x_{1}$, $x_{2}$) or constant values (e.g., $c_{1}$, $c_{2}$).
@@
Figure 2: Workflow for using LLMs for SR. First, the dataset is sent to LLMs (with or without context), which are instructed to suggest expressions without optimizing parameters. The generated expressions are then parsed by Python and optimized using SciPy (Nelder-Mead [77] method with basin-hopping as the numerical optimizer [97]). Results for each expression are stored in a Python dictionary and added to a list of results from prior expressions. The top-performing expressions are sent as feedback to LLMs, which are asked to suggest better expressions that optimize for both complexity and loss. The feedback loop is run for a set number of iterations.
Figure 3: Langmuir’s adsorption model illustrated in expression tree. The equation ‘complexity’ is computed from the equation string: each binary operator (+, -, *, /, **) is counted and weighted by 2, each occurrence of a unary function name (e.g. log, exp, sqrt, sin, etc.) is counted with weight 1, and a base cost of 1 is added so that even a constant expression scores at least 1.
@@
We use two prompts: 1) Initial Prompt - where we input data and ask GPT-4 and GPT-4o to suggest expressions, and 2) Iteration Prompt - where, along with data, we provide feedback in JSON format. We sort the expressions in descending order of mean squared error, then share this list as feedback, asking the LLMs to suggest new equations optimizing for complexity and loss. In addition to this, we also include a system message to guide the behavior of the language model, setting the tone of the conversation. We do not use chat history or any advanced forms of memory [7, 99, 100, 103] in this workflow. The history of generated expressions is maintained externally and provided as feedback in the iteration prompt. A Python function maintains this feedback loop by sorting and filtering a list of dictionaries based on MSE and complexity. At the start of the search, up to 6 expressions are always returned; later, the least accurate expressions that are not on the Pareto front are pruned in order to manage the length of the context window. By making each call to LLMs independent, we aim to minimize hallucinations that have been observed when the chat history becomes too large in conversational models [63, 88], as well as manage cost by pruning the large quantity of generated scratchpad text.
@@ 2.3 Prompt engineering
Removing bias: We aimed to make the workflow run smoothly without any human intervention, and therefore, it was important to obtain machine-readable and precise output from LLMs to ensure the SciPy function runs without any errors. A simple way to illustrate the expected outcome was by providing examples in a few-shot context [14, 91]. While this led to expressions matching the required syntax, we noticed the generated expressions resembled the examples we provided (Figure 4). While this taught the LLM correct syntax, it introduced bias that severely compromised the search.
Figure 4: Bias from prompt examples GPT-4 generated exact expressions from the prompt examples that were given to illustrate the output syntax. The revised prompts (Prompt 2 and 3) eliminate this bias from the equation prediction process.
This motivated our two-prompt setup, with an initial prompt tasking LLMs to generate unbiased expressions in LaTeX and an iteration prompt receiving previously-generated examples as Python script (now without bias). To bridge these, we converted the LaTeX text into SciPy-compatible text using a Python function for string formatting, which we developed after identifying the types of expressions GPT-4 was prone to generating. Although this approach did not completely resolve syntax errors, it effectively managed them for the GPT-4 model. While running the GPT-4o model, we realized that most syntax errors stemmed from incorrect LaTeX formatting. To address this, we made a minor adjustment by removing the term "LaTeX" and allowing the model to generate expressions as Python strings instead. This small change was sufficient to remove bias and was also less prone to syntax errors. We nervetheless processed the generated Python expressions through our formatting function, and it worked seamlessly.
@@
Figure 5: Effect of scratchpad on Langmuir Dataset. We observe substantial and qualitative improvements in the predicted expressions after implementing the scratchpad technique (Prompt 2). The suggested expressions for Langmuir’s dataset include operators (//) present in the target model ($y=\frac{c_{1}*x}{c_{2}+x})$.
@@
Consider scientific context: Our primary motivation for building this system was to test whether providing scientific context could shape the expressions generated by the LLM. SR programs often successfully generate expressions that fit the data well and are simple, yet they may not adhere to scientific principles or be otherwise “meaningful.” Yet scientists often have valuable insights into their domain that extend beyond these constraints, and they may not always know which specific expressions will best capture the nuances of their dataset or if entirely new expressions might be more effective. By incorporating scientific context, we aim to align and constrain the equation search to be consistent with scientific theories.
Classical SR approaches always incorporate some amount of guidance from the practitioners (e.g. by limiting available math operators and variables and incorporating some bias toward parsimonious expressions); more recent work enables SR programs to explicitly account for limiting behavior (shape-constrained SR) and dimensional constraints [85, 3, 29, 42, 68, 92, 12]. Incorporating these into SR algorithms typically requires bespoke modification of research software; our strategy is to use LLM prompting in natural language to instill this context, which might additionally include more fluid constraints like “generate diverse equations” and “consider scientific context.” Figure 21 in SI lists the context provided for each of the datasets.